Ocean waves are hitting a beach at a rate of 2 Hz. The distance between wave crests is 12 m. calculate the speed of the wave.
1
Expert's answer
2021-10-07T08:48:58-0400

"\lambda = 12m"

"f = 2Hz"

"v =\lambda*f = 12*2= 24\frac{m}{s}"

"\text{Answer: } 24\frac{m}{s}"